OpenStudy (anonymous):

Simplify the quotient 4.5x10^3 / 9x10^7 Write your answer in scientific notation. Show your work

OpenStudy (nikato):

What's 4.5/9?

OpenStudy (anonymous):

0.5

OpenStudy (nikato):

and remember dividing exponents with same base is just subtracting

OpenStudy (anonymous):

\[=4.5\times10^3\div9\times10^7\] \[=(4.5\div9)\times10^310^7\] =0.5*10^10 =?

OpenStudy (nikato):